# Noise-based reward-modulated learning

### v2

@@ 3 Results
3.1 Solving instantaneous and delayed reward problems
To ensure comparability with RMHL methods, which typically utilize single-hidden-layer networks, we first conduct experiments with a one-hidden-layer neural network for BP, NRL, and RMHL. Training details are provided in Appendix B. In Section titled “Scalability to multi-layer networks”, we extend the comparison to deeper networks with multiple hidden layers.
First, the Reaching problem [22], visualized in Fig. 1C, is an immediate reward problem that requires the system to reach and maintain a position at a target on a 1D ring by moving left, right, or remaining stationary. At each step, the system receives information about both the target’s position and its own, and a reward is provided based on proximity to the target over fixed-duration trials. Average performance across-trial and final performance (mean of the last 50 trials) are shown in Fig 1A and Fig 1B, respectively.
@@ 3.1 Solving instantaneous and delayed reward problems
Second, the Acrobot problem [53], visualized in Fig 1F, involves delayed reward and requires controlling a two-link robotic arm to reach a target height. At each time step, the system receives information about the angles and angular velocities of the two links and chooses one action: clockwise torque, counterclockwise torque, or no torque. Rewards are given based on the speed of completion, with a maximum time allowed. Average performance across trials and final performance (mean of the last 50 trials) are shown in Fig 1D and Fig 1E, respectively.
The third and most challenging problem, the Cartpole problem [1], visualized in Fig 1I, is a delayed reward problem where the system must balance a pole on a cart by moving left or right. At each time step, the system receives information about the cart’s position and velocity, along with the pole’s angle and angular velocity, and responds accordingly. Performance is measured by the time the pole remains balanced, with a maximum time allowed. Average performance across trials and final performance (mean of the last 50 trials) are shown in Fig 1G and Fig 1H, respectively. For all three tasks, NRL reaches the performance level of BP, whereas RMHL fails to do so.
3.2 Scalability to multi-layer networks
Here, we demonstrate that NRL can effectively assign credit in neural networks with multiple hidden layers; a challenging scenario where most biologically plausible algorithms struggle. We use both the Acrobot and the Cartpole tasks for this purpose, as they present a delayed reward problem, making them a more realistic test for credit assignment in reinforcement learning. In this comparison, we include the same baselines, BP and RMHL. Fig 2 displays the results for the Acrobot task and Fig 3 displays the results for the Cartpole tasks. The left panels of these two figures show performance across trials, averaged over 5 runs. In contrast, the right panels display the final performance (mean of the last 50 trials), also averaged over 5 runs. The results for neural networks with two hidden layers are shown in Fig 2A and Fig 2B, and 3A and B, and those for networks with three hidden layers are presented in Fig 2C and Fig 2D, and Fig 3C and Fig 3D.
@@ 3.2 Scalability to multi-layer networks
Our results indicate that NRL successfully learns to solve the tasks, achieving performance comparable to BP. In contrast, RMHL struggles with credit assignment in deeper networks. However, NRL requires more trials to converge as the network depth increases – an expected outcome due to the stochastic nature of the updates [24]. A similar trend is observed with BP, though to a lesser extent, as its gradient-based updates inherently provide more directed adjustments.
3.3 Using only noisy passes
In NRL, $\rho_{t}$ is calculated as the difference in the network’s output between the clean and noisy passes. However, a completely noiseless network is not feasible in more biophysically realistic scenarios. Instead, the clean network’s output can be approximated by averaging the outputs from multiple noisy passes, as the injected noise averages to zero $\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\xi_{t,i}^{l}=0$. This assumes that the network dynamics are faster than the environment dynamics, allowing the network to perform multiple forward passes before the environment changes.
To evaluate the accuracy of this approximation, we employ a single hidden-layer network and the Acrobot problem. The difference between the clean pass output and the averaged noisy passes output was calculated for each timestep and averaged over 500 timesteps, as shown in Fig 4A. Furthermore, Fig 4B illustrates the performance on the Acrobot problem using only noisy forward passes, starting with a minimum of 2 noisy passes. We also extended this evaluation to 10 noisy passes to explore the impact of increasing the number of passes, showing slightly faster initial convergence.
@@ 4 Discussion
In this work, we propose NRL, a noise-driven synaptic plasticity rule that bridges principles of reinforcement learning and optimization with biologically-inspired Hebbian updates. By leveraging stochastic neural activity, our method produces local synaptic updates modulated by a global reinforcement signal. It offers an efficient approach for neural adaptation, particularly suited to neuromorphic systems.
@@
We employ a top-down methodology, starting by deriving our learning rule from the mathematical principles of gradient-based optimization. We then approximate the gradients via directional derivatives and incorporate bottom-up constraints inspired by biological systems. Specifically, the gradient-approximation process is implemented through stochastic neural activity (aligning with the inherent noise in the brain [12, 51, 39]), and reinforcement learning mechanisms are realized through reward prediction errors (emulating dopaminergic reward signals [48, 47]), eligibility traces, and local synaptic plasticity. These elements enable learning under uncertainty [18]. This approach results in a learning rule that is both theoretically grounded and potentially realizable in neural circuits.
@@
Our results demonstrate the learning effectiveness of NRL, solving both immediate and delayed feedback/reward problems in simulated environments, achieving performance comparable to baselines optimized employing backpropagation. Additionally, it shows a significant improvement over RMHL [32, 40], a modern, well-known approach that employs noise as the core of its synaptic plasticity. Our scalability results further demonstrate that NRL effectively assigns credit even in multilayer networks, a challenge for many biologically plausible algorithms, indicating its potential for scaling. This is a critical difference with RMHL, which struggles to assign credit in multilayer networks evidenced by its minimal performance improvements across trials. Nevertheless, as the depth of the network increases, NRL shows slower convergence due to its noise-driven updates.
@@
NRL shares similarities with node perturbation methods [19, 24, 61, 58, 16], which inject noise into neurons to adjust weights based on state and performance fluctuations. However, unlike node perturbation – which requires constant feedback or reward for learning – NRL is well-suited for delayed reward setups. By leveraging eligibility traces [23, 26], it can effectively assign credit to past neural states even when actions and outcomes are separated in time.
@@
Gradient approximation, a cornerstone of our learning rule’s derivation, is also crucial in mathematical optimization and machine learning, enabling learning when exact gradients are computationally expensive or unavailable. Techniques such as forward gradient methods [2] and zeroth-order (ZO) optimization [35, 8] have emerged, utilizing stochastic perturbations and function value comparisons to estimate gradients. Recent advances, such as employing local auxiliary networks for informed gradient guesses [20], have significantly improved the alignment of these approximations with exact gradients, reducing the performance gap with backpropagation. These variance-reduction strategies could similarly enhance our approach by improving the quality of the updates and scalability to more complex architectures. Furthermore, compared to traditional gradient-based methods like backpropagation, gradient approximation methods offer a key advantage in computationally constrained scenarios, or when using non-differentiable networks, such as spiking neural networks [62, 56].
Additionally, the local-update structure of NRL provides compatibility with neuromorphic substrates [49], where efficient, low-power computations are essential. Neuromorphic systems can also inherently simulate noisy neural activity, enabling energy-efficient learning in applications where real-time adaptation is key. The inherent noise in NRL is also advantageous for reinforcement learning setups, as it naturally facilitates exploration-exploitation balance [27, 30]. Injected noise perturbs neural states, promoting exploratory deviations from expected values and thus introducing variations in the agent’s policy.
Our work has limitations that suggest future research directions. First, its dependence on multiple forward passes per synaptic update reduces biological plausibility, even though these passes could be performed in parallel. Future work should explore more efficient ways to estimate noise impacts, potentially eliminating the need for multiple passes. Second, our modeling approach deliberately involves a high degree of abstraction: we use rate-based neurons instead of spiking models, do not distinguish between excitatory and inhibitory populations, and treat synaptic potentiation and depression as a single continuum. While this provides clear insight into the capabilities and limitations of our approach, more biologically detailed models could offer further insights. Third, our reward prediction error calculation is a simplification. While sufficient for demonstrating the learning mechanism, the simplicity of using a running average may limit its effectiveness in scenarios where navigating through transient low-reward regions is necessary to achieve higher long-term rewards. In such cases, the model may underestimate the value of actions that lead to delayed benefits, potentially resulting in suboptimal updates. More sophisticated approaches, such as temporal difference learning [52, 54], or learned prediction through a critic network, could provide a richer and more temporally sensitive estimate of expected rewards. These strategies, which align more closely with biological dopaminergic signaling and can be implemented within actor-critic architectures, represent promising directions for improving learning performance in future work. Finally, while backpropagation may offer faster convergence in unconstrained settings, recent studies [9, 14], including ours, indicate that this gap can be narrowed for specific networks and tasks.
@@
Our findings highlight the potential of unifying optimization methods and neurobiological principles. By drawing inspiration from both domains, we have developed a learning rule that not only offers a biologically plausible alternative to backpropagation but also provides insights into how learning, particularly synaptic plasticity, might occur in biological neural systems. Conversely, the constraints of biological realism can guide the development of more efficient and robust artificial learning algorithms, particularly suited for neuromorphic hardware and real-time applications. This synergistic approach promises to advance our understanding of intelligence in both biological and artificial systems [64, 38], driving progress in neuroscience and machine learning alike.
@@ Appendix A Gradient approximation using directional derivatives
Now, substituting the expectation with a sampling under some empirical distribution,
@@
Finally, defining gradient descent on parameters $\theta$ as $\theta_{t+1}=\theta_{t}+\alpha\nabla g(\theta)$,

### v3

@@ 3 Results
3.1 Performance on control tasks with instantaneous and delayed rewards
@@
To ensure comparability with RMHL methods, which typically utilize single-hidden-layer networks, we first conduct experiments with a one-hidden-layer neural network for BP, NRL, and RMHL. Training details are provided in Appendix B. In Section titled “Scalability to deeper architectures”, we extend the comparison to deeper networks with multiple hidden layers.
First, the Reaching problem [21], visualized in Fig. 1C, is an immediate reward problem that requires the system to reach and maintain a position at a target on a 1D ring by moving left, right, or remaining stationary. At each step, the system receives information about both the target’s position and its own, and a reward is provided based on proximity to the target over fixed-duration trials. Average performance across-trial and final performance (mean of the last 50 trials) are shown in Fig. 1A and Fig. 1B, respectively.
Second, the Acrobot problem [55], visualized in Fig. 1F, involves delayed reward and requires controlling a two-link robotic arm to reach a target height. At each time step, the system receives information about the angles and angular velocities of the two links and chooses one action: clockwise torque, counterclockwise torque, or no torque. Rewards are given based on the speed of completion, with a maximum time allowed. Average performance across trials and final performance (mean of the last 50 trials) are shown in Fig. 1D and Fig. 1E, respectively.
The third and most challenging problem, the Cartpole problem [1], visualized in Fig. 1I, is a delayed reward problem where the system must balance a pole on a cart by moving left or right. At each time step, the system receives information about the cart’s position and velocity, along with the pole’s angle and angular velocity, and responds accordingly. Performance is measured by the time the pole remains balanced, with a maximum time allowed. Average performance across trials and final performance (mean of the last 50 trials) are shown in Fig. 1G and Fig. 1H, respectively. For all three tasks, NRL achieves a final performance comparable to the gradient-based baseline, BP, demonstrating a vast difference compared to RMHL.
3.2 Scalability to deeper architectures
Here, we demonstrate that NRL can effectively assign credit in neural networks with multiple hidden layers; a challenging scenario where most biologically plausible algorithms struggle. We use both the Acrobot and the Cartpole tasks for this purpose, as they present a delayed reward problem, making them a more realistic test for credit assignment in reinforcement learning. In this comparison, we include the same baselines, BP and RMHL. Figure 2 displays the results for the Acrobot task and Fig. 3 displays the results for the Cartpole tasks for neural networks consisting of two and three hidden layers. The left panels of these two figures show performance across trials, averaged over 5 runs. In contrast, the right panels display the final performance (mean of the last 50 trials), also averaged over 5 runs.
@@ 3.2 Scalability to deeper architectures
Our results indicate that NRL successfully learns to solve the tasks, achieving performance comparable to BP. In contrast, RMHL struggles with credit assignment in deeper networks. However, NRL requires more trials to converge as the network depth increases, which is an expected outcome due to the stochastic nature of the updates [23]. A similar trend is observed with BP, though to a lesser extent, as its gradient-based updates inherently provide more directed adjustments.
3.3 Learning using only noisy passes
In NRL, $\rho_{t}$ is calculated as the difference in the network’s output between the clean and noisy passes. However, generating a perfectly noiseless pass may be hardware-infeasible or energy-prohibitive in neuromorphic circuits. Instead, we show that the clean network’s output can be approximated by averaging the outputs from multiple noisy passes, as the injected noise averages to zero $\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\xi_{t,i}^{l}=0$. This assumes that the network dynamics are faster than the environment dynamics, allowing the network to perform multiple forward passes before the environment changes.
To evaluate the accuracy of this approximation, we employ a single hidden-layer network and the Acrobot problem. The difference between the clean pass output and the averaged noisy passes output was calculated for each timestep and averaged over 500 timesteps, as shown in Fig. 4A. Furthermore, Fig. 4B illustrates the performance on the Acrobot problem using only noisy forward passes, starting with a minimum of 2 noisy passes. We also extended this evaluation to 10 noisy passes to explore the impact of increasing the number of passes, showing slightly faster initial convergence.
@@ 4 Discussion
In this work, we propose NRL, noise-based reward-modulated learning, a novel synaptic plasticity rule that bridges the principles of reinforcement learning and optimization with biologically inspired Hebbian updates. By leveraging stochastic neural activity, our method produces local synaptic updates modulated by a global reinforcement signal. It offers an efficient approach for neural adaptation, particularly suited to neuromorphic systems.
@@
We employ a top-down methodology, deriving our learning rule from the mathematical principles of gradient-based optimization. We approximate the gradients through directional derivatives and incorporate bottom-up constraints inspired by biological systems. Specifically, the gradient-approximation process is implemented through stochastic neural activity, directly aligning with the inherent noise in neuromorphic substrates [50, 45, 29] and the brain [11, 53, 37]. Reinforcement learning mechanisms are realized through reward prediction errors (emulating dopaminergic reward signals [49, 48]), eligibility traces, and local synaptic plasticity. These elements also enable learning under uncertainty [17]. This approach results in a learning rule that is both theoretically grounded and potentially realizable in event-driven neuromorphic hardware.
@@
Our results demonstrate the learning effectiveness of NRL, solving both immediate and delayed feedback/reward problems in simulated environments, achieving performance comparable to baselines optimized employing backpropagation. In addition, it shows a significant improvement over RMHL [33, 38], a modern approach that also uses noise as the core of its synaptic plasticity. Our scalability results further demonstrate that NRL effectively assigns credit even in multilayer networks, a challenge for many biologically plausible or local algorithms, indicating its potential for scaling. This is a critical difference from RMHL, which struggles to assign credit in multilayer networks, as evidenced by its minimal performance improvements across trials. Nevertheless, as the depth of the network increases, NRL shows slower convergence due to its noise-driven updates.
@@
NRL shares similarities with node perturbation methods [18, 23, 65, 62, 15], which inject noise into neurons to adjust weights based on state and performance fluctuations. However, unlike node perturbation, which requires constant feedback or reward for learning, NRL is well-suited for delayed reward setups. By leveraging eligibility traces [22, 25], it can effectively assign credit to past neural states even when actions and outcomes are separated in time.
@@
Gradient approximation, a cornerstone of our learning rule’s derivation, is also crucial in mathematical optimization and machine learning, enabling learning when exact gradients are computationally expensive or unavailable. Techniques such as forward gradient methods [2] and zeroth-order optimization [34, 6] have emerged, utilizing stochastic perturbations and function value comparisons to estimate gradients. Recent advances, such as employing local auxiliary networks for informed gradient guesses [19], have significantly improved the alignment of these approximations with exact gradients, reducing the performance gap with backpropagation. These variance-reduction strategies could similarly enhance our approach by improving the quality of the updates and scalability to more complex architectures. Furthermore, compared to traditional gradient-based methods like backpropagation, gradient approximation methods offer a key advantage in computationally constrained scenarios, or when using non-differentiable networks, such as spiking neural networks [66, 58].
Additionally, the local-update structure of NRL provides direct compatibility with neuromorphic substrates [50], where only locally available information in the synapse is prioritized and global communication is severely penalized. This eliminates the need for the high-bandwidth global data movement and large memory buffers required by backpropagation, the primary source of energy consumption in deep learning models, addressing the fundamental energy and latency bottlenecks of conventional Von Neumann systems. Furthermore, NRL’s reliance on noise for gradient approximation offers a paradigm shift: it transforms the device stochasticity common in neuromorphic circuits from an engineering challenge into a computational resource. The inherent noise in NRL is also advantageous for reinforcement learning setups, as it naturally facilitates the exploration-exploitation balance [26, 31]. Injected noise perturbs neural states, promoting exploratory deviations from expected values and thus introducing variations in the agent’s policy, which benefits adaptation on neuromorphic edge devices.
Despite its promise, our work has limitations that suggest future research directions. The current requirement for multiple forward passes (noisy and clean) per synaptic update, while theoretically simple, introduces a time-latency overhead. Future work should explore more efficient ways to estimate noise impacts, potentially eliminating the need for multiple passes. In addition, while we focused on a rate-based model, the deployment of NRL on physical hardware like Intel Loihi [8] requires its extension to spiking neural networks (SNNs). Furthermore, the simplicity of our RPE calculation limits its robustness in more complex tasks requiring long-term planning. Future development should incorporate more sophisticated temporal difference learning mechanisms [54, 56], such as a learned value function within an actor-critic architecture [28], to provide a richer temporally sensitive estimate of rewards. Finally, while backpropagation may offer faster convergence in unconstrained settings, recent studies [7, 13], including ours, indicate that this gap can be narrowed for specific networks and tasks.
@@
Our findings ultimately highlight the potential for unifying optimization methods and neurobiological principles. By drawing inspiration from both domains, we have developed a learning rule that not only offers an alternative to backpropagation but is also algorithmically suited for low-energy, event-driven neuromorphic hardware. In the same way, the constraints of biological realism can guide the development of more efficient and robust artificial learning algorithms. This synergistic approach provides a crucial new path for creating low-energy, fast, and adaptive AI systems [68, 36].
@@ Appendix A Gradient approximation using directional derivatives
Now, substituting the expectation with a sampling under some empirical distribution:
@@
Finally, defining gradient descent on parameters $\theta$ as $\theta_{t+1}=\theta_{t}+\alpha\nabla g(\theta)$: